two pounds of 44.4% potassium bromate solution at 70 degrees celsius has been cooled at 10 degrees celsius. Estimate the amount of crystals to be recovered
1
Expert's answer
2020-10-20T08:36:59-0400

Solubility of KBr at 10"C^\circ" is 60 g of KBr /100 g water

mass percent of KBr in this solution is

"w(KBr) = \frac{mass (KBr)}{mass (solution)}=\frac{60}{60+100}=0.375"

1) Find mass of KBr in the given solution

"2 pounds (solution)\times \frac{453.592 g}{1 pound}= 907.185 g (solution)"

"mass(KBr) = 907.185 g \times 0.444 =402.790 g"


2) x- mass of the crystalls

mass percent of KBr in the solution after the crystals precipitated

"mass \% = \frac{mass(KBr)}{mass(solution)}=\frac{402.790-x}{907.185-x}"

what is equal to mass percent of KBr at 10 "C^\circ" ", 0.375"

solve the equation

"0.375=\frac{402.790-x}{907.185-x}"

x"=100.16"

the mass of the crystals is 100.16 g